OpenStudy (anonymous):

Explain how to find the angle of inclination of a line if you know the slope of the line.

OpenStudy (anonymous):

gogo satellite!

OpenStudy (anonymous):

it is \[\tan^{-1}(\theta)\]

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

maybe?

OpenStudy (anonymous):

oh wait. sorry. i confused myself and wrote \[\theta\] which is silly.

OpenStudy (anonymous):

\[\tan \alpha=\Delta y/\Delta x=slope\]Solve for angle of inclination,\[\alpha=\tan^{-1} (slope)\]

OpenStudy (anonymous):

it is inverse tan, you guys are good.

OpenStudy (anonymous):

yes of course. it is \[\tan^{-1}(m)\]

OpenStudy (anonymous):

or course no one would call the slope theta, but i guess you could...

OpenStudy (anonymous):

The slope itself isn't theta. Tan (theta) would be the slope :)

OpenStudy (anonymous):

THANKS GUYS!